OpenStudy (anonymous):

An eight sided dice has its faces with 2,4,6,8...16. what is the expected result of a normal roll? The choices are 7,8,9 and 16

OpenStudy (mathstudent55):

\(\large E(x) = x_1p_1 + x_2p_2 + ... x_np_n\) Here, \(x_1 = 2, x_2 = 4, ..., x_8 = 16\) Also, \(p_1 = \dfrac{1}{8}, p_2 = \dfrac{1}{8}, ..., x_8 = \dfrac{1}{8}\)

OpenStudy (mathstudent55):

That means all you have to do is add all values and divide by 8.